A hiker begins a 25 km(point B) hike from point A at 7am travelling at a rate of 4km per hour. At 9:30, a bike rider leaves point B travelling at a rate of 16 km per hour. What time will the hiker and biker cross paths (meet)? At what kilometer will they meet?
Answer by mananth(16949)

Hiker 4 km/h
Biker 16 km/h 2.5 hours late
Distance 25 miles
In 2.5 hours Hiker has traveled 10 miles before Biker starts
So distance = 15 when Biker starts
Let time they meet be x hours
Hiker distance traveled 4x miles
Biker distance traveled 16(15 -x)

Time taken to meet is same
t=d/r
combined speed= 20 mph

20x=15
x= 0.75 hours
Hiker 4x=3 miles
Hiker would have traveled 3+10=13 miles
Biker (25-13)=12 miles
They will cross at 10:15 AM